# Consistent Supervised-Unsupervised Alignment for Generalized Category Discovery

### v2

@@ 1 Introduction
Generalized Category Discovery (GCD) has raised emerging attention in recent years, which aims to simultaneously classify known categories and discover novel ones from unlabeled data. Unlike traditional semi-supervised learning, which assumes all categories are predefined during training, GCD operates in a more realistic open-world scenario zhang2021openworld ; bendale2015openset ; liu2020openvocabulary where only a small portion of known-category data is labeled, and all novel-category samples lack both labels and category information. It requires models to utilize limited supervision from labeled known categories while autonomously discovering semantically coherent clusters in the unlabeled subset without assuming predefined category structures or a clear distinction between known and novel distributions.
Recent research in GCD has witnessed substantial progress. Early research GCD2022 introduces a contrastive learning framework to distinguish known and novel categories. On this basis, contrastive learning-based methods DCCL2023 ; PromptCAL2023 refine category boundaries using dynamic contrastive learning and prompt-based affinity learning. Additionally, representation learning-based methodsPIM2023 ; CMS2024 ; AGCD2024 ; SelEx2024 focus on optimizing feature representations to enhance category separability.
@@
Neural Collapse (NC) is a phenomenon in well-trained classification networks, where features of each category align with a Simplex Equiangular Tight Frame (ETF) structure Papyan2020_NeuralCollapse . In this structure, within-category features collapse to their respective category mean, and the means of different categories are at the vertices of a simplex. This alignment maximizes inter-category separation while maintaining within-category compactness, creating an optimal geometric arrangement for classification tasks.
@@ 2.1 Generalized Category Discovery (GCD)
GCD identifies known and novel categories within a partially labeled dataset. Early methodsGCD2022 leveraged contrastive learning to distinguish categories. Later, DCCL DCCL2023 and PromptCAL PromptCAL2023 refined category boundaries using dynamic contrastive learning and prompt-based affinity mechanisms. GPC GPC2023 and SimGCD SimGCD2023 introduced Gaussian Mixture Models and parametric classification frameworks to estimate category distributions. Other methods have focused on self-supervised clustering and feature refinement. PIM PIM2023 and CMS CMS2024 introduced contrastive mean-shift learning and representation alignment strategies to improve novel category separability. Reciprocal learning and class-wise regularization have been introduced in RLCD Liu2025RLCD , while ProtoGCD Ma2025ProtoGCD unifies prototype optimization and debiasing for category discovery. Meanwhile, approaches like UNO Fini2021_UNO and ORCA Cao2022_ORCA explored semi-supervised representation learning, while RankStats Han2020_RankStats employed ranking-based statistics to facilitate unknown category discovery. Meanwhile, there is a growing line of work on continual generalized category discovery Kim2023CGCD ; Cendra2024PromptCCD ; Park2024OCGCD ; Rypesc2024CAMP . However, most existing methods rely on dynamically learned prototypes and lack a consistent optimization target. This results in category confusion and imbalanced learning.
@@ 2.2 Neural Collapse and its Applications
Neural Collapse (NC) describes a geometric phenomenon in deep networks where, in the final stage of training, category feature representations converge to form a Simplex Equiangular Tight Frame structure Papyan2020_NeuralCollapse . This structure ensures optimal category separation and minimal intra-category variance, making it highly effective for classification. Several works have provided theoretical insights into Neural Collapse under different loss functions, demonstrating that it emerges naturally in both cross-entropy loss Graf2021_DissectingContrastive ; Fang2021_LayerPeeled ; Ji2022_Unconstrained and mean squared error (MSE) loss Mixon2020_UnconstrainedNC ; Poggio2020_ExplicitReg ; Zhou2022_ForwardFSCIL ; Han2022_NeuralCollapseMSE . Beyond theoretical exploration, NC has been applied to various machine-learning tasks. In imbalanced learning, NC principles have been leveraged to stabilize classifier prototypes under category-imbalanced training Yang2022b_LearnableClassifier ; Zhong2023_ImbalancedSegmentation . In few-shot class-incremental learning, studies have explored ETF-based classifiers to enhance feature-classifier alignment and mitigate forgetting yang2023neural . In federated learning, NC has been shown to improve model generalization under non-iid settings Huang2023_NCFederatedLearning . Similar prototype-based mechanisms have also been applied in domain-incremental scenarios, where consistent concept representations are maintained across domains Wang2025DualCP . Furthermore, NC-inspired approaches have been developed to address large-scale classification problems, such as optimizing one-vs-rest margins for generalized NC Jiang2023_GeneralizedNC . TRAILER Xiao2024TRA also draws inspiration from NC; however, it employs a fixed classifier with a cross-entropy–based ETF loss applied to both labeled and pseudo-labeled data, which can cause optimization bias and instability. Our NC-GCD framework extends NC principles by pre-assigning ETF prototypes to known and novel categories. By integrating supervised and unsupervised ETF alignment, our method effectively addresses GCD’s misalignment and category confusion.
@@ 3.2 Neural Collapse and Definition of Simplex ETF
Neural Collapse, observed during the terminal training phase in well-regularized neural networks on balanced datasets Papyan2020_NeuralCollapse , describes a symmetric geometric structure in the last-layer features and classifier weights. Features of the same category collapse to their within-category mean, which aligns with the corresponding classifier weights, forming a Simplex ETF.
@@ 4.2 ETF Alignment
To align features with the pre-assigned ETF prototype $\mathbf{P}=\{p_{1},p_{2},\dots,p_{K}\}$, we select the top $\alpha\%$ of samples with the highest similarity $s_{i}$ within each cluster. These high-confidence samples are pulled toward the corresponding ETF prototype $p_{k}$ using a Dot-Regression LossYang2022b_LearnableClassifier :
@@ 4.4 Comparative Representation Learning
Following CMSCMS2024 , we adopt the unsupervised loss encourages similarity between embeddings of the same sample while distinguishing them from other samples:
@@ 5.1 Experimental Setup
Datasets. Our evaluation spans six image classification benchmarks, including both generic and fine-grained datasets. For generic datasets, we use CIFAR-100 krizhevsky2009learning and ImageNet-100 deng2009imagenet ; tian2020contrastive . For fine-grained datasets, we evaluate on CUB-200 wah2011caltech , Stanford Cars krause20133d , FGVC Aircraft maji2013fine , and Herbarium19 tan2019herbarium . To separate categories into known and novel categories, we follow the SSB split protocol li2020ssb for the fine-grained datasets. For CIFAR-100 and ImageNet-100, we perform a random category split using a fixed seed, consistent with previous studies.
Implementation Details. Our method utilizes the pre-trained DINO ViT-B/16 caron2021emerging ; dosovitskiy2021image as the backbone network. We fine-tune the final layer of the image encoder along with a projection head, following prior work GCD2022 ; CMS2024 . The projection head consists of an MLP with a 768-dimensional input, a 2048-dimensional hidden layer, and a 768-dimensional output, followed by a GeLU activation function hendrycks2016gaussian . The learning rate is set to 0.1. Other hyperparameters, including batch size, temperature $\tau_{s}$, weight decay $\lambda$, and the number of augmentations, are set to 128, 0.07, $1e^{-4}$, and 2, respectively, in accordance with previous studies. All experiments are conducted on an NVIDIA 3090 GPU. Further implementation details can be found in the Appendix.
@@ 5.2 Comparison with State of the Art Methods
We evaluate our NC-GCD framework on several fine-grained and generic classification benchmarks to demonstrate its efficacy in GCD. Table 1 presents a comparison between our method and the previous methods GCD2022 ; DCCL2023 ; PromptCAL2023 ; GPC2023 ; SimGCD2023 ; PIM2023 ; CMS2024 ; SPT2024 ; SelEx2024 ; Xiao2024TRA , evaluated on both fine-grained and generic datasets with and without access to the ground-truth (GT) number of categories K for clustering. For the case without access to GT K, we estimate the K through clustering, which is then used to align the ETF prototypes. Meanwhile, to comprehensively evaluate the performance of various methods across Fine-Grained Datasets, Generic Classification Datasets, and All datasets, Table 2 presents the average performance.
@@ A.1 Training Parameters and Details
In this study, we evaluate the NC-GCD framework on various datasets using a pre-trained DINO ViT-B/16 model as the backbone. The final layer of the image encoder is fine-tuned alongside a projection head for both supervised and unsupervised alignment tasks, as per previous works GCD2022 ; CMS2024 . The projection head is a Multi-Layer Perceptron (MLP) consisting of a 768-dimensional input layer, a 2048-dimensional hidden layer, and a 768-dimensional output layer, followed by a GeLU activation function hendrycks2016gaussian . We empirically set periodic clustering epoch T=5 as a trade-off between stability and computational efficiency. Larger T may reduce adaptation speed, while smaller T increases computational cost.
The learning rate is set to 0.1, and other key hyperparameters such as batch size (128), temperature ($\tau_{s}=0.07$), weight decay ($1e^{-4}$), and the number of augmentations (2) follow the settings used in prior worksGCD2022 . All experiments are conducted using an NVIDIA 3090 GPU, which provides sufficient power for the fine-grained operations required in our method.
@@ A.2 Dataset Details and Category Splits
CIFAR100 krizhevsky2009learning 80 20
ImageNet100 deng2009imagenet 50 50
CUB-200-2011 wah2011caltech 100 100
Stanford Cars krause20133d 98 98
FGVC-Aircraft maji2013fine 50 50
Herbarium19 tan2019herbarium 341 342
CIFAR-100. CIFAR-100 krizhevsky2009learning consists of 60,000 images across 100 categories. For the GCD evaluation, we used 80 categories as known categories for training and the remaining 20 as unknown categories for testing.
ImageNet100. ImageNet100 deng2009imagenet is a subset of ImageNet with 100 categories. Following standard protocols, we randomly assign 50 categories as known categories and the other 50 as unknown categories.
CUB-200-2011. CUB-200-2011 wah2011caltech is a fine-grained bird dataset with 200 categories. We split the dataset into 100 known and 100 unknown categories following the standard fine-grained GCD protocol li2020ssb .
Stanford Cars. Stanford Cars krause20133d is a fine-grained dataset with 196 car categories. We use 98 categories for training (known categories) and 98 for testing (unknown categories), consistent with previous fine-grained GCD work.
FGVC Aircraft. FGVC Aircraft maji2013fine includes 100 aircraft categories. The dataset is split into 50 known and 50 unknown categories, adhering to the SSB protocol.
Herbarium19. Herbarium19 tan2019herbarium is a large-scale fine-grained dataset with 683 plant species. Due to its high intra-category variance, we split the dataset into 341 known and 342 unknown categories to evaluate GCD on more challenging fine-grained datasets.
Category Split Protocol. For general datasets (CIFAR-100 and ImageNet100), we apply a random category split using a fixed seed to ensure reproducibility. For fine-grained datasets (CUB-200-2011, Stanford Cars, and FGVC Aircraft), we use the SSB split protocol li2020ssb , which ensures balanced category distributions while preserving intra-category variance. For Herbarium19, the dataset is split into nearly equal known and unknown categories, testing the model’s robustness in handling datasets with significant intra-category variation.
@@ B.5 Performance under Estimated Category Number $K$
We conduct an initial CMS-style CMS2024 representation learning phase for 30 epochs and estimate the number of categories every 5 epochs based on the current feature space. Following our earlier analysis on the impact of ETF prototype number deviations, we adopt a conservative estimation strategy by selecting the maximum estimated $K$ across all estimation points. This approach effectively reduces the negative impact of underestimating $K$, which, as previously shown, leads to significant performance degradation on novel categories, while the model remains more tolerant to overestimation.
@@ D.2 Training Process
Fig. 8 illustrates the evolution of feature representations for all categories across different training epochs. At epoch 0, the features are highly entangled, with significant category confusion due to the lack of geometric constraints. As training progresses, the feature clusters gradually become more distinct, aligning with their respective ETF prototypes. By epoch 40, the initial separation between categories emerges, and by epoch 100, the majority of categories are well-structured. By epoch 200, the features exhibit a near-complete collapse to a simplex ETF structure, aligning closely with the Neural Collapse (NC) theory Papyan2020_NeuralCollapse , where features within each class concentrate around their mean while inter-class distances remain maximized.
@@ Appendix F Broader Impacts
This work explores a novel framework for Generalized Category Discovery (GCD), enhancing the ability of AI systems to autonomously recognize and differentiate new concepts in open-world environments. By reducing the dependence on large-scale labeled datasets, our method lowers the barriers to applying AI technologies in domains where data annotation is expensive or difficult to obtainGCD2022 ; han2025learn .
The proposed framework encourages more efficient use of unlabeled data, offering practical solutions for knowledge discovery in fields such as environmental monitoring, healthcare diagnostics, and industrial quality control. Furthermore, it has the potential to accelerate the development of adaptive intelligent systems capable of continuously learning from new, unseen data without extensive human supervision. We hope this research will contribute to broader advancements in learning with limited supervision and inspire further innovations in building more flexible and scalable AI systemsGong2024Preface .

### v3

@@ 1 Introduction
Generalized Category Discovery (GCD) has raised emerging attention in recent years, which aims to simultaneously classify known categories and discover novel ones from unlabeled data. Unlike traditional semi-supervised learning, which assumes all categories are predefined during training, GCD operates in a more realistic open-world scenario Li et al. (2025); Bendale and Boult (2015); Joseph et al. (2021) where only a small portion of known-category data is labeled, and all novel-category samples lack both labels and category information. It requires models to utilize limited supervision from labeled known categories while autonomously discovering semantically coherent clusters in the unlabeled subset without assuming predefined category structures or a clear distinction between known and novel distributions.
Recent research in GCD has witnessed substantial progress. Early research Vaze et al. (2022) introduces a contrastive learning framework to distinguish known and novel categories. On this basis, contrastive learning-based methods Pu et al. (2023); Zhang et al. (2023) refine category boundaries using dynamic contrastive learning and prompt-based affinity learning. Additionally, representation learning-based methodsChiaroni et al. (2023); Choi et al. (2024); Ma et al. (2024); Rastegar et al. (2024) focus on optimizing feature representations to enhance category separability.
@@
Neural Collapse (NC) is a phenomenon in well-trained classification networks, where features of each category align with a Simplex Equiangular Tight Frame (ETF) structure Papyan et al. (2020). In this structure, within-category features collapse to their respective category mean, and the means of different categories are at the vertices of a simplex. This alignment maximizes inter-category separation while maintaining within-category compactness, creating an optimal geometric arrangement for classification tasks.
@@ 2.1 Generalized Category Discovery (GCD)
GCD identifies known and novel categories within a partially labeled dataset. Early methodsVaze et al. (2022) leveraged contrastive learning to distinguish categories. Later, DCCL Pu et al. (2023) and PromptCAL Zhang et al. (2023) refined category boundaries using dynamic contrastive learning and prompt-based affinity mechanisms. GPC Zhao et al. (2023) and SimGCD Wen et al. (2023) introduced Gaussian Mixture Models and parametric classification frameworks to estimate category distributions. Other methods have focused on self-supervised clustering and feature refinement. PIM Chiaroni et al. (2023) and CMS Choi et al. (2024) introduced contrastive mean-shift learning and representation alignment strategies to improve novel category separability. Reciprocal learning and class-wise regularization have been introduced in RLCD Liu et al. (2025), while ProtoGCD Ma et al. (2025) unifies prototype optimization and debiasing for category discovery. Meanwhile, approaches like UNO Fini et al. (2021) and ORCA Cao et al. (2021) explored semi-supervised representation learning, while RankStats Han et al. (2020) employed ranking-based statistics to facilitate unknown category discovery. Meanwhile, there is a growing line of work on continual generalized category discovery Kim et al. (2023); Cendra et al. (2024); Park et al. (2024); Rypeść et al. (2024). However, most existing methods rely on dynamically learned prototypes and lack a consistent optimization target. This results in category confusion and imbalanced learning.
@@ 2.2 Neural Collapse and its Applications
Neural Collapse (NC) describes a geometric phenomenon in deep networks where, in the final stage of training, category feature representations converge to form a Simplex Equiangular Tight Frame structure Papyan et al. (2020). This structure ensures optimal category separation and minimal intra-category variance, making it highly effective for classification. Several works have provided theoretical insights into Neural Collapse under different loss functions, demonstrating that it emerges naturally in both cross-entropy loss Graf et al. (2021); Fang et al. (2021); Ji et al. (2021) and mean squared error (MSE) loss Mixon et al. (2022); Poggio and Liao (2020); Zhou et al. (2022); Han et al. (2021). Beyond theoretical exploration, NC has been applied to various machine-learning tasks. In imbalanced learning, NC principles have been leveraged to stabilize classifier prototypes under category-imbalanced training Yang et al. (2022); Zhong et al. (2023). In few-shot class-incremental learning, studies have explored ETF-based classifiers to enhance feature-classifier alignment and mitigate forgetting Yang et al. (2023). In federated learning, NC has been shown to improve model generalization under non-iid settings Huang et al. (2023). Similar prototype-based mechanisms have also been applied in domain-incremental scenarios, where consistent concept representations are maintained across domains Wang et al. (2025). Furthermore, NC-inspired approaches have been developed to address large-scale classification problems, such as optimizing one-vs-rest margins for generalized NC Jiang et al. (2023). TRAILER Xiao et al. (2024) also draws inspiration from NC; however, it employs a fixed classifier with a cross-entropy–based ETF loss applied to both labeled and pseudo-labeled data, which can cause optimization bias and instability. Our NC-GCD framework extends NC principles by pre-assigning ETF prototypes to known and novel categories. By integrating supervised and unsupervised ETF alignment, our method effectively addresses GCD’s misalignment and category confusion.
@@ 3.2 Neural Collapse and Definition of Simplex ETF
Neural Collapse, observed during the terminal training phase in well-regularized neural networks on balanced datasets Papyan et al. (2020), describes a symmetric geometric structure in the last-layer features and classifier weights. Features of the same category collapse to their within-category mean, which aligns with the corresponding classifier weights, forming a Simplex ETF.
@@ 4.2 ETF Alignment
To align features with the pre-assigned ETF prototype $\mathbf{P}=\{p_{1},p_{2},\dots,p_{K}\}$, we select the top $\alpha\%$ of samples with the highest similarity $s_{i}$ within each cluster. These high-confidence samples are pulled toward the corresponding ETF prototype $p_{k}$ using a Dot-Regression LossYang et al. (2022):
@@ 4.4 Comparative Representation Learning
Following CMSChoi et al. (2024), we adopt the unsupervised loss encourages similarity between embeddings of the same sample while distinguishing them from other samples:
@@ 5.1 Experimental Setup
Datasets. Our evaluation spans six image classification benchmarks, including both generic and fine-grained datasets. For generic datasets, we use CIFAR-100 Krizhevsky and Hinton (2009) and ImageNet-100 Deng et al. (2009). For fine-grained datasets, we evaluate on CUB-200 Wah et al. (2011), Stanford Cars Krause et al. (2013), FGVC Aircraft Maji et al. (2013), and Herbarium19 Tan et al. (2019). To separate categories into known and novel categories, we follow the SSB split protocol Vaze et al. (2021) for the fine-grained datasets. For CIFAR-100 and ImageNet-100, we perform a random category split using a fixed seed, consistent with previous studies.
Implementation Details. Our method utilizes the pre-trained DINO ViT-B/16 Caron et al. (2021); Dosovitskiy (2020) as the backbone network. We fine-tune the final layer of the image encoder along with a projection head, following prior work Vaze et al. (2022); Choi et al. (2024). The projection head consists of an MLP with a 768-dimensional input, a 2048-dimensional hidden layer, and a 768-dimensional output, followed by a GeLU activation function Hendrycks and Gimpel (2016). The learning rate is set to 0.1. Other hyperparameters, including batch size, temperature $\tau_{s}$, weight decay $\lambda$, and the number of augmentations, are set to 128, 0.07, $1e^{-4}$, and 2, respectively, in accordance with previous studies. All experiments are conducted on an NVIDIA 3090 GPU. Further implementation details can be found in the Appendix.
@@ 5.2 Comparison with State of the Art Methods
We evaluate our NC-GCD framework on several fine-grained and generic classification benchmarks to demonstrate its efficacy in GCD. Table 1 presents a comparison between our method and the previous methods Vaze et al. (2022); Pu et al. (2023); Zhang et al. (2023); Zhao et al. (2023); Wen et al. (2023); Chiaroni et al. (2023); Choi et al. (2024); Wang et al. (2024); Rastegar et al. (2024); Xiao et al. (2024), evaluated on both fine-grained and generic datasets with and without access to the ground-truth (GT) number of categories K for clustering. For the case without access to GT K, we estimate the K through clustering, which is then used to align the ETF prototypes. Meanwhile, to comprehensively evaluate the performance of various methods across Fine-Grained Datasets, Generic Classification Datasets, and All datasets, Table 2 presents the average performance.
@@ 5.3 Ablation Study
Ablation study of ETF loss coefficient $\beta$ , more results of the effectiveness of ETF Alignment and Parameter analysis will be presented in the Appendix.
@@ A.1 Training Parameters and Details
In this study, we evaluate the NC-GCD framework on various datasets using a pre-trained DINO ViT-B/16 model as the backbone. The final layer of the image encoder is fine-tuned alongside a projection head for both supervised and unsupervised alignment tasks, as per previous works Vaze et al. (2022); Choi et al. (2024). The projection head is a Multi-Layer Perceptron (MLP) consisting of a 768-dimensional input layer, a 2048-dimensional hidden layer, and a 768-dimensional output layer, followed by a GeLU activation function Hendrycks and Gimpel (2016). We empirically set periodic clustering epoch T=5 as a trade-off between stability and computational efficiency. Larger T may reduce adaptation speed, while smaller T increases computational cost.
The learning rate is set to 0.1, and other key hyperparameters such as batch size (128), temperature ($\tau_{s}=0.07$), weight decay ($1e^{-4}$), and the number of augmentations (2) follow the settings used in prior worksVaze et al. (2022). All experiments are conducted using an NVIDIA 3090 GPU, which provides sufficient power for the fine-grained operations required in our method.
@@ A.2 Dataset Details and Category Splits
CIFAR100 Krizhevsky and Hinton (2009) 80 20
ImageNet100 Deng et al. (2009) 50 50
CUB-200-2011 Wah et al. (2011) 100 100
Stanford Cars Krause et al. (2013) 98 98
FGVC-Aircraft Maji et al. (2013) 50 50
Herbarium19 Tan et al. (2019) 341 342
CIFAR-100. CIFAR-100 Krizhevsky and Hinton (2009) consists of 60,000 images across 100 categories. For the GCD evaluation, we used 80 categories as known categories for training and the remaining 20 as unknown categories for testing.
ImageNet100. ImageNet100 Deng et al. (2009) is a subset of ImageNet with 100 categories. Following standard protocols, we randomly assign 50 categories as known categories and the other 50 as unknown categories.
CUB-200-2011. CUB-200-2011 Wah et al. (2011) is a fine-grained bird dataset with 200 categories. We split the dataset into 100 known and 100 unknown categories following the standard fine-grained GCD protocol Vaze et al. (2021).
Stanford Cars. Stanford Cars Krause et al. (2013) is a fine-grained dataset with 196 car categories. We use 98 categories for training (known categories) and 98 for testing (unknown categories), consistent with previous fine-grained GCD work.
FGVC Aircraft. FGVC Aircraft Maji et al. (2013) includes 100 aircraft categories. The dataset is split into 50 known and 50 unknown categories, adhering to the SSB protocol.
Herbarium19. Herbarium19 Tan et al. (2019) is a large-scale fine-grained dataset with 683 plant species. Due to its high intra-category variance, we split the dataset into 341 known and 342 unknown categories to evaluate GCD on more challenging fine-grained datasets.
Category Split Protocol. For general datasets (CIFAR-100 and ImageNet100), we apply a random category split using a fixed seed to ensure reproducibility. For fine-grained datasets (CUB-200-2011, Stanford Cars, and FGVC Aircraft), we use the SSB split protocol Vaze et al. (2021), which ensures balanced category distributions while preserving intra-category variance. For Herbarium19, the dataset is split into nearly equal known and unknown categories, testing the model’s robustness in handling datasets with significant intra-category variation.
@@ B.5 Performance under Estimated Category Number $K$
We conduct an initial CMS-style Choi et al. (2024) representation learning phase for 30 epochs and estimate the number of categories every 5 epochs based on the current feature space. Following our earlier analysis on the impact of ETF prototype number deviations, we adopt a conservative estimation strategy by selecting the maximum estimated $K$ across all estimation points. This approach effectively reduces the negative impact of underestimating $K$, which, as previously shown, leads to significant performance degradation on novel categories, while the model remains more tolerant to overestimation.
@@ D.2 Training Process
Fig. 8 illustrates the evolution of feature representations for all categories across different training epochs. At epoch 0, the features are highly entangled, with significant category confusion due to the lack of geometric constraints. As training progresses, the feature clusters gradually become more distinct, aligning with their respective ETF prototypes. By epoch 40, the initial separation between categories emerges, and by epoch 100, the majority of categories are well-structured. By epoch 200, the features exhibit a near-complete collapse to a simplex ETF structure, aligning closely with the Neural Collapse (NC) theory Papyan et al. (2020), where features within each class concentrate around their mean while inter-class distances remain maximized.
@@ Appendix F Broader Impacts
This work explores a novel framework for Generalized Category Discovery (GCD), enhancing the ability of AI systems to autonomously recognize and differentiate new concepts in open-world environments. By reducing the dependence on large-scale labeled datasets, our method lowers the barriers to applying AI technologies in domains where data annotation is expensive or difficult to obtainVaze et al. (2022); Han et al. (2025).
The proposed framework encourages more efficient use of unlabeled data, offering practical solutions for knowledge discovery in fields such as environmental monitoring, healthcare diagnostics, and industrial quality control. Furthermore, it has the potential to accelerate the development of adaptive intelligent systems capable of continuously learning from new, unseen data without extensive human supervision. We hope this research will contribute to broader advancements in learning with limited supervision and inspire further innovations in building more flexible and scalable AI systemsGong and Wang (2024).